OpenStudy (anonymous):

Evaluate u/z + xy^2 for u = 20, x = 4, y = 7, z = 10. A. 786 B. 198 C. 294 D. 900

OpenStudy (anonymous):

@Nnesha Do you think you can solve some more of them? :)

OpenStudy (anonymous):

and also, do I have to put every question in different posts, or can I put them all here? :)

Nnesha (nnesha):

substitute u for 20 x for 4 y for 7 and z for 10 then solve

Nnesha (nnesha):

well the rule is one question per post so other user will able to help you as wll

OpenStudy (anonymous):

so...20/10 + 4 times 7^2?

Nnesha (nnesha):

yep right!

OpenStudy (anonymous):

that would mean...198? Just making sure :)

Nnesha (nnesha):

hm how ?

OpenStudy (anonymous):

(20/10) 2 + 4 x 7^2 = 198

OpenStudy (anonymous):

Just making sure :)

Nnesha (nnesha):

\[\frac{ 20 }{ 10+4(7)^2 }\]

OpenStudy (anonymous):

20/10 = 2, 2 + 4(7)^2 = 198....thats how I got it.

Nnesha (nnesha):

wait \[\frac{ u }{ z } +xy^2\] is this ur question ?

OpenStudy (anonymous):

Yep

Nnesha (nnesha):

ohhh okay then 198 is right i thought it's \[\frac{ x }{ z+xy^2 }\]

OpenStudy (anonymous):

Thanks then lol

Nnesha (nnesha):

np